# FootstepNet: an Efficient Actor-Critic Method for Fast On-line Bipedal Footstep Planning and Forecasting

### v1

@@ I Introduction
The goal of footstep planning is to find a suitable sequence of footsteps to achieve the desired task. Some work focuses on finding proved-to-be-feasible sequences, which is difficult in complex and cluttered environments. In this work, we only consider 2D environment with one obstacle. We consider the problem of finding the minimal sequence. Even in this simplified form, this optimization problem has no closed-form solution and is challenging to compute efficiently on-line. Additionally, navigating in substantially more complex environments can be achieved with a global path planner, using intermediate targets for the local planner as presented in [6] and in figure 2.
@@
Finally, reinforcement learning (RL) approaches were used. Hofer and Rouxel [17] proposed a RL-based approach to produce walk orders to approach a soccer ball. However, they did not consider obstacles. Meduri and al. proposed DeepQ stepper [18], a footstep planning based on DQN algorithm [19], but their approach was exclusively focused on preserving the robot stability while tracking a provided velocity task. In DeepGait [20] by Tsounis et al., quadrupedal locomotion is separated in two DRL subproblems: the Gait Planner (GP) and Gait Controller. GP addresses a geometrical problem which is the quadruped equivalent of the footstep planning. Even if feasibility is extensively checked at this stage, it is unclear that the reward function encourages a minimization of the number of steps because of its complexity.
@@
The second contribution is FootstepNet forecasting : the method’s ability to provide an estimated number of footsteps required to reach different candidates of local targets. Those estimations do not depend on computations of all the required footsteps and can, therefore, produce quick and useful insights for upstream decision-making (see figure 1).
@@ II Problem Statement
We assume the footstep planning to be part of a broader system, as depicted in figure 2. In this context, an upstream autonomous decision is made to select the target footstep. To do so, a global overview of the environment can be used (e.g., using the shortest path with A*-like methods). In our approach, like in [6], we do not consider finding all the footsteps to reach a distant target. We rather focus on navigating efficiently in the vicinity of the robot. Because of the dynamic nature of the environment, as well as slippages and perturbations, replanning become inevitable and jeopardizes long-term plans. Moreover, long-range navigation is more likely to asymptotically comply with simple heuristics, such as assuming a constant velocity. However, with the minimization of the number of footsteps in mind, local navigation can yield complex maneuvers as presented in figure 3
Figure 3: Example of footsteps generated by FootstepNet planning for the three possible goals of figure 1 – The target positions are close to each other, however the generated footsteps to reach them use different complex maneuvers.
@@ IV Method
A footstep displacement $\Delta\phi=(\Delta x,\Delta y,\Delta\theta)$, is parametrized as on figure 4. It describes the pose of the swing foot in the frame of the support foot. When a support swap occurs, the swing foot becomes the new support foot, producing a new footstep. A sequence of footsteps can then be built from successive displacements, which defines the trajectory.
Figure 5: Overview of the proposed method – First, offline training is carried out during which the agent learns the policy by interacting with the simulated geometric environment. During online inference, we then use the trained networks to, on the one hand, estimate the number of steps using the critic and, on the other hand, to determine the sequence of steps to be performed using the actor.
@@
Given a target $\phi_{t}=(f_{t},x_{y},y_{t},\theta_{t})$, the goal of the footstep planning problem is to find a sequence $\Phi_{p}=(\phi_{r},\phi_{2},\dots,\phi_{t})$ such that displacements are feasible, and with minimal length $|\Phi_{p}|$. This problem is non-linear because of the possible rotations of the robot. It also has non-convex constraints because of the obstacle avoidance, but also possibly because of the shape of $\mathcal{F}$. For those reasons, there are no known closed-form solutions.
@@ IV-2 Action-space
An action $a\in\mathcal{A}=\mathbb{R}^{3}$ is a tuple: $a=\Delta\phi$, as specified in figure 4. The actions are clipped to lie in the feasible set $\mathcal{F}$. After applying back the symmetry operator $\sigma$ on $\Delta y$ and $\Delta\theta$, such a displacement can be integrated to obtain a new footstep.
@@ V Experiments
Using the proposed method described in Sec. IV, an agent dedicated to the footsteps planning was trained. The computed policy network is used to generate footsteps to reach a given target –FootstepNet planner– and the critic network to forecast the number of steps to the same aim –FootstepNet forecast– (cf. figure 5) The main objectives of our experiments were thus to evaluate networks’ planning and forecasting performances and to demonstrate the feasibility of the whole pipeline during a real-world scenario : the RoboCup Competition.
@@ V-B2 On-board inference
The first experiment is dedicated to compare the performances of FootstepNet planner against the state-of-the-art footstep planner ARA* [16] [15]. We used implementation made by J. Garimort and al. [22] in the ROS footstep planner package [^2] https://wiki.ros.org/footstep_planner. In order to do so, three different scenarios were created (cf. figure 6) :
@@ V-C Footstep forecasting
The second experiment aims to validate the accuracy of FootstepNet forecast compared to the roll-out of the policy. The forecasting predicts the number of steps to reach the target from a given state. The roll-out is the number of steps generated by the FootstepNet planning until the target is reached. The same scenarios as in the first experiment were used (cf. figure 7) with the addition of four different obstacle sizes for the AO and FO situations. However, for each randomly selected starting pose, the critic was compared to the roll-out on a set of choices between multiple close/near equivalent targets. Table II shows that the forecasting is nearly as effective as the roll-out to select the best target to reach among these near equivalent ones (eg. For facing a 0.15m obstacle, the critic-based erroneous ratio is only 0.59%). According to the low mean estimation relative error (4.55% for avoiding a 0.15m obstacle), we can deduce that FootstepNet forecast is also able to accurately predict the number of steps to reach a target.
@@ V-D Deployment on a kid-size humanoid during RoboCup 2023
The robots are custom-made and fully autonomous, carrying their battery and computational power. For the 2023 edition, we deployed FootstepNet in Sigmaban, competing in the kid-size category. In a competition setup, it is crucial to take as few footsteps as possible to be as fast as possible. In particular, many maneuvers are necessary around the ball to get in position to kick (see figure 1). Let us mention that this task, i.e. reach a position as quickly as possible by walking, is natural beyond the context of soccer, and would be useful in many other applications. Let us also mention that the competition context requires a high level of reliability. The footsteps were re-planned periodically with an horizon of $H=5$ footsteps, yielding a constant computation time of $225\mu s$. Footsteps are then passed to the downstream whole-body planner to plan the CoM trajectory with a scheme similar to [30]. To decide the target kick and placement, an upstream strategy/decision-making module was designed. Relevant information (position of the ball, allies and opponents) was used to select target poses, using an estimated time-to-goal score. Since several poses were often equivalent, FootstepNet forecasting was then used to make the final choice (figure 9).
@@
Additionally, we provide a complementary video [^3] https://youtu.be/EL1rJh45vug about FootstepNet and demonstrates its application on a Sigmaban robot.

### v2

@@ I Introduction
The goal of footstep planning is to find a suitable sequence of footsteps to achieve the desired task. Some work focuses on finding proved-to-be-feasible sequences, which is difficult in complex and cluttered environments. In this work, we only consider 2D environment with one obstacle. We consider the problem of finding the minimal sequence. Even in this simplified form, this optimization problem has no closed-form solution and is challenging to compute efficiently on-line. Additionally, navigating in substantially more complex environments can be achieved with a global path planner, using intermediate targets for the local planner as presented in [6] and in Fig. 2.
@@
Finally, reinforcement learning (RL) approaches were used. Hofer and Rouxel [17] proposed a RL-based approach to produce walk orders to approach a soccer ball. However, they did not consider obstacles. Meduri and al. proposed DeepQ stepper [18], a footstep planning based on DQN algorithm [19], but their approach was exclusively focused on preserving the robot stability while tracking a provided velocity task. Güldenstein [20] proposed a footstep planning method that was compared to an expert method but does not focus on minimizing the number of steps. In DeepGait [21] by Tsounis et al., quadrupedal locomotion is separated in two DRL subproblems: the Gait Planner (GP) and Gait Controller. GP addresses a geometrical problem which is the quadruped equivalent of the footstep planning. Even if feasibility is extensively checked at this stage, it is unclear that the reward function encourages a minimization of the number of steps because of its complexity.
@@
The second contribution is FootstepNet forecasting : the method’s ability to provide an estimated number of footsteps required to reach different candidates of local targets. Those estimations do not depend on computations of all the required footsteps and can, therefore, produce quick and useful insights for upstream decision-making (see Fig. 1).
@@ II Problem Statement
We assume the footstep planning to be part of a broader system, as depicted in Fig. 2. In this context, an upstream autonomous decision is made to select the target footstep. To do so, a global overview of the environment can be used (e.g., using the shortest path with A*-like methods). In our approach, like in [6], we do not consider finding all the footsteps to reach a distant target. We rather focus on navigating efficiently in the vicinity of the robot. Because of the dynamic nature of the environment, as well as slippages and perturbations, replanning become inevitable and jeopardizes long-term plans. Moreover, long-range navigation is more likely to asymptotically comply with simple heuristics, such as assuming a constant velocity. However, with the minimization of the number of footsteps in mind, local navigation can yield complex maneuvers as presented in Fig. 3
Figure 3: Example of footsteps generated by FootstepNet planning for the three possible goals of Fig. 1 – The target positions are close to each other, however the generated footsteps to reach them use different complex maneuvers.
@@ IV Method
A footstep displacement $\Delta\phi=(\Delta x,\Delta y,\Delta\theta)$, is parametrized as on Fig. 4. It describes the pose of the swing foot in the frame of the support foot. When a support swap occurs, the swing foot becomes the new support foot, producing a new footstep. A sequence of footsteps can then be built from successive displacements, which defines the trajectory.
Figure 5: Overview of the proposed method – First, offline training is carried out during which the agent learns the policy by interacting with the RL environment. During online inference, we then use the trained networks to, on the one hand, estimate the number of steps using the critic and, on the other hand, to determine the sequence of steps to be performed using the actor.
@@
Given a target $\phi_{t}=(f_{t},x_{t},y_{t},\theta_{t})$, the goal of the footstep planning problem is to find a sequence $\Phi_{p}=(\phi_{r},\phi_{2},\dots,\phi_{t})$ such that displacements are feasible, and with minimal length $|\Phi_{p}|$. This problem is non-linear because of the possible rotations of the robot. It also has non-convex constraints because of the obstacle avoidance, but also possibly because of the shape of $\mathcal{F}$. For those reasons, there are no known closed-form solutions.
@@ IV-2 Action-space
An action $a\in\mathcal{A}=\mathbb{R}^{3}$ is a tuple: $a=\Delta\phi$, as specified in Fig. 4. The actions are clipped to lie in the feasible set $\mathcal{F}$. After applying back the symmetry operator $\sigma$ on $\Delta y$ and $\Delta\theta$, such a displacement can be integrated to obtain a new footstep.
@@ V Experiments
Using the proposed method described in Sec. IV, an agent dedicated to the footsteps planning was trained. The computed policy network is used to generate footsteps to reach a given target –FootstepNet planner– and the critic network to forecast the number of steps to the same aim –FootstepNet forecast– (cf. Fig. 5) The main objectives of our experiments were thus to evaluate networks’ planning and forecasting performances and to demonstrate the feasibility of the whole pipeline during a real-world scenario : the RoboCup Competition.
@@ V-B2 On-board inference
The first experiment is dedicated to compare the performances of FootstepNet planner against the state-of-the-art footstep planner ARA* [16] [15]. We used implementation made by J. Garimort and al. [23] in the ROS footstep planner package [^2] https://wiki.ros.org/footstep_planner. In order to do so, three different scenarios were created (cf. Fig. 6) :
@@ V-C Footstep forecasting
The second experiment aims to validate the accuracy of FootstepNet forecast compared to the roll-out of the policy. The forecasting predicts the number of steps to reach the target from a given state. The roll-out is the number of steps generated by the FootstepNet planning until the target is reached. The same scenarios as in the first experiment were used (cf. Fig. 7) with the addition of four different obstacle sizes for the AO and FO situations. However, for each randomly selected starting pose, the critic was compared to the roll-out on a set of choices between multiple close/near equivalent targets. Table II shows that the forecasting is nearly as effective as the roll-out to select the best target to reach among these near equivalent ones (eg. For facing a 0.15m obstacle, the critic-based erroneous ratio is only 0.59%). According to the low mean estimation relative error (4.55% for avoiding a 0.15m obstacle), we can deduce that FootstepNet forecast is also able to accurately predict the number of steps to reach a target.
@@ V-D Deployment on a kid-size humanoid during RoboCup 2023
The robots are custom-made and fully autonomous, carrying their battery and computational power. For the 2023 edition, we deployed FootstepNet in Sigmaban, competing in the kid-size category. In a competition setup, it is crucial to take as few footsteps as possible to be as fast as possible. In particular, many maneuvers are necessary around the ball to get in position to kick (see Fig. 1). Let us mention that this task, i.e. reach a position as quickly as possible by walking, is natural beyond the context of soccer, and would be useful in many other applications. Let us also mention that the competition context requires a high level of reliability. The footsteps were re-planned periodically with an horizon of $H=5$ footsteps, yielding a constant computation time of $225\mu s$. Footsteps are then passed to the downstream whole-body planner to plan the CoM trajectory with a scheme similar to [31]. To decide the target kick and placement, an upstream strategy/decision-making module was designed. Relevant information (position of the ball, allies and opponents) was used to select target poses, using an estimated time-to-goal score. Since several poses were often equivalent, FootstepNet forecasting was then used to make the final choice (Fig. 9).
@@
Additionally, we provide a video [^3] https://youtu.be/sLu-LHdO6Mk about FootstepNet and demonstrates its application on a Sigmaban robot.
@@ VI Conclusion
The open-source code for this project is available on our GitHub repository [^4] https://github.com/Rhoban/footstepnet_envs for further use and contributions.